OpenStudy (anonymous):

will medal and fan solve -7/3 n=-42

OpenStudy (mvpriest6):

how do you start this

OpenStudy (mvpriest6):

do u know

OpenStudy (anonymous):

not at all

OpenStudy (mvpriest6):

oh okay hold on then

OpenStudy (mvpriest6):

what is 7/3

OpenStudy (mvpriest6):

@iGreen

OpenStudy (anonymous):

-7/3 it a fraction

OpenStudy (mvpriest6):

2.333333333333333

OpenStudy (mvpriest6):



OpenStudy (mvpriest6):

r u sure this is a real question

OpenStudy (mvpriest6):

@briggs

OpenStudy (mvpriest6):

what grade r u in

OpenStudy (anonymous):

yes this is the question

OpenStudy (anonymous):

7th

OpenStudy (mvpriest6):

cause this is hard

OpenStudy (anonymous):

ikr

OpenStudy (mvpriest6):

@iGreen plz help

OpenStudy (mvpriest6):

brb

OpenStudy (anonymous):

N=\[\frac{ 1 }{ 8 }\]

OpenStudy (igreen):

Okay, so we have: \(- \dfrac{7}{3}n = -42\) Now we have to divide -7/3 to both sides, or multiply by its reciprocal: \(n = -\dfrac{42}{1} \times -\dfrac{3}{7}\) Can you multiply those? @briggs

OpenStudy (anonymous):

yea hold on a sec

OpenStudy (anonymous):

133/3 or 44 1/3

OpenStudy (anonymous):

is that right

OpenStudy (igreen):

No..

OpenStudy (igreen):

\(-\dfrac{42}{1} \times -\dfrac{3}{7} \rightarrow \dfrac{-42 \times -3}{-1 \times -7}\) Multiply..what's -42 * -3?

OpenStudy (anonymous):

126

OpenStudy (igreen):

Yes, and what's -1 * -7?

OpenStudy (anonymous):

7

OpenStudy (igreen):

Yes, that gives us: \(\dfrac{126}{7}\) Can you divide that?

OpenStudy (anonymous):

18

OpenStudy (igreen):

Yep, that's your answer.